OpenStudy (anonymous):

A 4.00% NaOH solution by mass has a density of 1.054 g/mL. What is the minimum molarity of an HCl(aq) solution that can be used to titrate a 8.00 mL sample of the NaOH(aq) if the titration is to be accomplished without having to refill a 50.00 mL buret used in the titration?

OpenStudy (anonymous):

I'm not sure why they're giving you the % by mass. I'm thinking that you can convert the 1.054g to mole and convert the mL to L and then you'll have molarity.

OpenStudy (anonymous):

Once you have molarity you can figure out the number of moles in 8.00ml

OpenStudy (anonymous):

Yea, that confused me too, I'll try doing it without that first.

OpenStudy (anonymous):

I'm not 100% sure but I think they threw it in to complicate the question.

OpenStudy (anonymous):

Yea, that might be.

OpenStudy (anonymous):

How do I take the molarity and use it to find moles in 8 mL? Because I have 26.35 as the molarity.

OpenStudy (anonymous):

But 4% means that the solution is only 4% NaOH probably, so maybe i do 4% of the molarity of what i have now?

OpenStudy (anonymous):

Give me a few minutes. I'm in middle of trying to figure it out.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

how did you get molarity to be 26.35? 1.054g/39.997g/m = 0.02635 mol. But you need to divide it by volume. I'll show you how to do it in one step. It's kind of long but you solve the mol of 8.00 ml in one step and it will make your teacher happy because it's using dimensional analysis. If it doesn't make any sense. I'll try to do it step wise.

OpenStudy (anonymous):

Ok, the molarity was because I had 39.997/1.054 instead of the other way around.

OpenStudy (anonymous):

I can't do it in one step. So I'll break it up. \[\frac{ 1.054gNaOH }{ 1 }*\frac{ molNaOH }{ 39.997gNaOH } = 0.02635 molNaOH\] 1 mL = 0.001L \[\frac{ 0.02635molNaOH }{ 0.001L }= 26.35M\] Your calculation seems correct. I'm just concerned because M is usually not this high. But we should be able to verify by using the %mass Now you should be able to convert 8.00mL to mole

OpenStudy (anonymous):

How would I do that? 8 x (mol/mL)?

OpenStudy (anonymous):

\[\frac{ mL }{ 1 }*\frac{ L }{ mL }*\frac{ mole }{ L }\]

OpenStudy (anonymous):

Ok, that makes sense. Basically it is mL x (mol/mL) because L is 1 anyway, but you do have to put it in anyway, so ok.

OpenStudy (anonymous):

I'm sorry, I made a mistake the density given 1.054g/mL is for the entire solution so molarity can't be calculated from it.

OpenStudy (anonymous):

If we do 4% of that density could it work?

OpenStudy (anonymous):

I think so. What if we multiply the whole thing by 1,000, so we'll have 1,054g/1000mL which is 1,054g/L 4% of 1,054 is 0.04(1,054) that will be the mass of NaOH. If we convert it to mole we'll have the molarity. What do you think?

OpenStudy (anonymous):

Yea, I was trying something like that. So far my work looks something like: 1.054 x (1000/1) x (1/40) x (4/100) = 1.054 M NaOH, could that work?

OpenStudy (anonymous):

Why do you have the 1/40?

OpenStudy (anonymous):

I'm not sure to be honest, I was just thinking that too...I'm trying to find where that's from.

OpenStudy (anonymous):

I think without it it looks good.

OpenStudy (anonymous):

What's weird is that our molarity from our previous answer was all that minus the 4% But 4% of our previous answer is my answer

OpenStudy (anonymous):

Without the 1/40 it doesn't come out as 4% of the previous answer...

OpenStudy (anonymous):

OOOoh 1/40 is the mass of NaOH, it's to get to moles.

OpenStudy (anonymous):

Correct I was going to do it as a separate step, but you can do it in one step.

OpenStudy (anonymous):

Ok, so after that, I did (8/1) x (1/1000) x (1.054/1) = 0.008432, is that right?

OpenStudy (anonymous):

Besides sig figs, at least.

OpenStudy (anonymous):

The molarity comes out to be the same as the density.

OpenStudy (anonymous):

Oh yeah, it does. That's not good.

OpenStudy (anonymous):

0.00432 is correct. I think it is good, I can't prove it, but the 4% and 1/40 from the molar mass somehow gives us the original number, if the molar mass would be something else we would get a different number.

OpenStudy (anonymous):

I made a typo 0.008432 is correct

OpenStudy (anonymous):

Ah, ok. How many sig figs? 3?

OpenStudy (anonymous):

3, but your not done yet. 0.008432 is the number of moles you need to figure out the molarity of HCL

OpenStudy (anonymous):

So I convert 0.00843 to HCl?

OpenStudy (anonymous):

Yes, using the 50mL from the buret

OpenStudy (anonymous):

I'm lost on how I would do that, 50 x what?

OpenStudy (anonymous):

50 ml * 1/1000 would convert it to L To figure out molarity I would not use dimensional analysis (I'm not sure it's even possible) M = mole / volume solvent. So divide the number of moles by the the volume of the buret

OpenStudy (anonymous):

0.00843 as moles or HCl's formula mass?

OpenStudy (anonymous):

Because M = HCl's mass per mole / L buret holds is WAAAY too high.

OpenStudy (anonymous):

0.00843 was the number of moles NaCl in 8.00mL NaCl and HCl react ina 1:1 ratio We do not need formulas for HCl. Molarity is moles / volume (If the buret would be 20 ml the HCl would need to be more concentrated) Since the buret is 50mL it can be less concentrated, so the volume can be 50ml. Now just plug the numbers in to calculate the molarity.

OpenStudy (anonymous):

I mis-typed; NaCl and HCl react in a 1:1 ratio. The formula mass for HCl is not necessary.

OpenStudy (anonymous):

Ok, so it's just 0.00843/0.05?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

That was right! Yay!

OpenStudy (anonymous):

Great! Sorry I went off track at first, but glad you got it. Does it make sense what we're doing?

OpenStudy (anonymous):

Yea, for the most part. Some parts are still a bit hazy, but it's a bit more clear now. Anyway, go to sleep now, don't let me cut off your sleep any further.

OpenStudy (anonymous):

Good night

OpenStudy (anonymous):

Good night